OpenStudy (anonymous):

Convert the given from of the function into the other form. 1. y=(x+2)(x-3) into standard and vertex form vertex form y-k = a(x-h)^2

OpenStudy (anonymous):

@swissgirl do you know how to do this?

OpenStudy (anonymous):

standard form requires you to multiply out

OpenStudy (anonymous):

\[(x+2)(x-3)=x^2-3x+2x-6=x^2-x+6\]

OpenStudy (anonymous):

I need help with vertex form @satellite73 I understand standard but not vertex.

OpenStudy (saifoo.khan):

Vertex is the turning point.